10x^3+15x^2+12x+18=0 i am having trouble working it cant seem to solve

Answer by scott8148(6628)

grouping ___ (10x^3 + 15x^2) + (12x + 18) = 0

factoring ___ (2x + 3)(5x^2) + (2x + 3)(6) = 0

rearranging ___ (2x + 3)(5x^2 + 6) = 0